OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -12, -7, -2, ... an = -17 + 5(n + 2) an = -17 + 5(n + 1) an = -17 + 5(n - 1) an = -17 x 5(n - 1)

OpenStudy (anonymous):

@dpaInc @phi @hartnn

hartnn (hartnn):

could u find the common difference ?

hartnn (hartnn):

or u know what is common difference ?

OpenStudy (anonymous):

no i need help with everything :/

hartnn (hartnn):

common difference = next term - current term so what is -12 - (-17) = ? -7 - (-12) = ? and so o n

OpenStudy (anonymous):

5

hartnn (hartnn):

so thats your d= common difference = 5 1st term = a = -17 and the general formula for n'th term = ? u know ?

OpenStudy (anonymous):

no :(

hartnn (hartnn):

Tn = a+(n-1)d now just put the values

OpenStudy (anonymous):

im confused :(

hartnn (hartnn):

d=5, a= -17 any problems ?

OpenStudy (anonymous):

and n?

hartnn (hartnn):

keep n as n only

hartnn (hartnn):

u have n in options

OpenStudy (anonymous):

tn=-17+(n-1)5

hartnn (hartnn):

that is correct

OpenStudy (anonymous):

oh thats the answer?

hartnn (hartnn):

yes, that the final answer, don't u see it in options ?

OpenStudy (anonymous):

the last one.

hartnn (hartnn):

no,

hartnn (hartnn):

the last one has multiplication sign, x

OpenStudy (anonymous):

the one above the last one

hartnn (hartnn):

yeah

OpenStudy (anonymous):

ty

hartnn (hartnn):

welcome ^_ ^